Thursday 11.06.25 Jenn/Gus



Manager Communication: Stacked team tonight. Couple of events. Good coverage

Customer Comments: Some Southern folks came by today, was cool to get some support from reps who just want to be here.

BAR Staff Observations: Solid crew tonight. Matthew and Natasha crushed it. Spoons was a breeze

FOH Staff Observations: Great teamwork from the floor tonight

Hiring/Training: Reached out to a recent former employee today. Hoping to hear back from them tomorrow.

Sick Leave/ Call outs:

Special Events: Spoons. Screening in Th.3, brought some peeps to the bar.

Staff Meeting Agenda notes:

86ed / counts on items:

Prep List: On the board for Tevon. Jenn made a bunch o stuff.

R&D comp checks:

Over/Short on Banks:

Pay Outs: 
$150.00 (Halloween Costume Prize) JH

Lost & Found?:

Ordering/Receiving/Shipping: 
Amazon (bar cleaning supplies + coffee lids) + Baldor + Empire + Harney + Manhattan + M.S. Walker + Sarene + Terrestrial + To Rise + Union arrived. JH

Maintenance: It was slow tonight so I took it upon myself to change out the first thermal sensor in the upstairs. It was odd. I replaced it as normal, but afterwards there were things that gave me pause. Lets see how this machine does. I filled it with water and left it on Cold 3. If in the morning it is simply cold we know we are on the right track. I am not going to go into my fears at this moment because as a great man once said, "Now is not the time for fear, that comes later." Regardless, I think we started the fire.

Mobilebytes Issues:

Filmbot Issues:

Comments

Popular posts from this blog

Sunday 07.06.2025 Ian/Lauren

Friday 7/4/25 Gus/Lauren/Ariel NPP Bar

Sunday 4/6/25 Ian/Lauren